# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

### v3

@@ 8. Proof of the Main Theorem
Let us apply the construction of Section 5. This construction works modulo Claims 5.2, 5.4, 5.6, 5.7. Claims 5.2 and 5.6 follow from Lemma 7.1. Claims 5.4 and 5.7 follow from Lemma 7.5. The union $\bigcup\limits_{i=1}^{n}M_{i}$ consist always only of $(1,d-1)$-fans, thus the final set of edges $M$ is a $(1,d-1)$-matching.
@@
For every $u\in U$ there is a step where an edge incident to $u$ is added to $M$. Then if the copy $v_{u}$ was not added to $M$ earlier, in the second part of this step $v_{u}$ is also added to $M$. It follows that $M$ is a perfect $(1,d-1)$-matching of the graph $\Gamma$. This property also implies computability of this matching. Indeed, the construction guarantees that for any $n$ all edges from $M$ of the forms $(n,j)$ and $(k,n)$ belong to $\bigcup\limits_{i=1}^{n}M_{i}$. Therefore there is an algorithm which computes all such pairs from $M$ for any $n$. We see that the function $f$ realized by this matching is computable.
@@
It remains to show that $f$ has controlled sizes of its cycle. Since the edges $(u_{0},v^{0}_{0,1})$ and $(v_{u_{0}},u_{v^{0}_{0,1}})$ are added to $M$ at step 1 we see $f^{2}(u_{0})=f(u_{v^{0}_{0,1}})=u_{0}$. Since $u_{0}=1$, condition $(i)$ of Definition 2.4 is satisfied.
@@
Until the end of the proof we will abuse the notation and use a natural number $n$ both for vertex and number of step. Note that as a vertex $n$ appears in $M$ at the $n$-th step at latest. This follows from minimality of $u_{n}$ in $U^{(n-1)}$. The length of a cycle created at the $n$-th step cannot be greater than $\max\{2,n\}$ so if $n$ is in a cycle then $f^{i}(n)=n$ for some $i\leq n$ and condition $(ii)$ of Definition 2.4 is satisfied.
@@
It remains to show that condition $(iii)$ is satisfied. Let $\mathfrak{f}_{i}$ be the partial function (living in $\mathbb{N}$) realized by $M_{i}$ and $f_{n}$ be the partial function realized by $\bigcup\limits_{i=0}^{n}M_{i}$. Let $F_{n}=\{(f_{n}(m),m):m\in\mathsf{Dom}(f_{n})\}$ be the graph of $f_{n}$ on $\mathsf{Dom}(f_{n})\cup\mathsf{Rng}(f_{n})$. Then $M_{n}=\{(\mathfrak{f}_{n}(m),m):m\in\mathsf{Dom}(\mathfrak{f}_{n})\}$ is the graph of $\mathfrak{f}_{n}$ and is a subgraph of $F_{n}$ too.
@@
Observe that $F_{n}$ has at most $n+1$ connected components. Indeed, each subgraph $M_{i}$ has exactly one connected component. Furthermore, one of the following properties holds:
  • $M_{n}$ is a connected component of $F_{n}$;
  • $M_{n}$ is a fan with a root which already appears in $F_{n-1}$ as a vertex of degree $1$ (see the way how fans of $U$-elements are added to $M$).
@@
The construction guarantees that $F_{n}$ consists only of vertices of degree $1$ and $d$. When a vertex has degree $1$, then its $F_{n}$-neighbourhood is exactly its $f_{n}$-image. When a vertex has degree $d$, then its $f_{n}$-image and $d-1$ preimages are in $F_{n}$. In particular each connected component of $F_{n}$ contains a cycle. The length of the cycle is not greater than $\max\{n,2\}$.
@@
Since the value $f_{n-1}(n)$ is defined, $n$ belongs to some connected component of $f_{n-1}$. Thus there exist $k$ and $l$ such that $f_{n-1}^{k+l}(n)=f_{n-1}^{k}(n)$. These $k$ and $l$ work for the equality $f^{k+l}(n)=f^{k}(n)$. It remains to show that $k$ can be bounded by $2n$ and $l$ by $n$.
@@
The latter estimate is easy: the biggest cycle that can be constructed before the $n$-th step does not have more than $\max\{n-1,2\}$ elements. Below we will bound $l$ by $n$ for simplicity.
@@
In order to show that $k$ is bounded by $2n$ let us estimate the size of a subset of $U$ without a cycle that can be added to the matching $M$ in the process of the $n$-th step of the construction. It must consist of elements of $U^{(n-1)\perp}$ added to the matching at the iteration of part $2$ of the $n$-th step together with $u_{n}$. Therefore we can bound it by the maximal possible number of elements in $U^{(n-1)\perp}$ increased by $1$.
@@
Let us denote the number of elements from $U^{(s)\perp}$ added to $M$ at the $s$-th step by $\ell_{s}$. Clearly, $f^{\ell_{s}+1}(n)\in M_{j}$ for some $s\leq n$ and $j\leq n-1$. If no cycle is constructed in the $j$-th step then $f^{\ell_{j}+1}(f^{\ell_{s}+1}(n))\in M_{i}$ for some $i\leq j-1$. Iterating this argument we arrive at
@@
Since at the $n$-th step the size of $\bigcup\limits_{s=1}^{n}U^{(s)\perp}$ does not exceed $n-1$,
@@ 9. Schneider’s theorem and computable entourages of coarse spaces
Let $d\geq 3$. Let a coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that for any finite $F\subseteq\mathbb{N}$ we have $|E[F]|\geq(d+2)|F|$. Then there exists a computable $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a $d$-regular forest. Moreover, there exist an algorithm which for any $m,n\in\mathbb{N}$ recognizes if $m$ and $n$ are in the same connected component of $\Gamma(E^{\prime})$.
@@
Let $E$ be a highly computable symmetric entourage as in the formulation. Consider the graph $\Gamma(R)$ defined for the symmetric relation $R:=E\setminus\Delta_{\mathbb{N}}\subseteq\mathbb{N}\times\mathbb{N}$. Since the coarse space $(\mathbb{N},\mathcal{E})$ is a bounded geometry, the neighbourhood of each vertex is finite. The graph $\Gamma(R)$ is highly computable by high computability of $E$.
We are going to apply Theorem 4.1 to $\Gamma(R)$. It remains to show that $\Gamma(R)$ satisfies $c.e.H.h.c.(d)$. To see this note:
@@
In particular the inequality $n\leq|X|$ implies $n\leq|N(X)|-d|X|\leq|N(X)|-\frac{1}{d}|X|$. Since the identity map on $\mathbb{N}$ is a total computable function, it follows that $\Gamma(R)$ satisfies $c.e.H.h.c.(d)$.
Let us apply Theorem 4.1. We obtain a computable function $f$ realizing a perfect $(1,d-1)$ matching in the graph $\Gamma(R)$ with controlled sizes of its cycles. In order to construct $E^{\prime}\in\mathcal{E}$ as in the formulation we adapt the proof of Theorem 2.2 of [Sch18].
Since $f$ is a total, computable, surjective, $(d-1)$ to $1$ function, the graph of $f$ (denoted by $\Gamma(f)$) is computable and $d$-regular. We remind the reader that $f$ satisfies the following properties:
@@
$P(f)=\{n\in\mathbb{N}\,|\,\exists m\geq 1(f^{m}(n)=n)\}\,\mbox{ and }\,P_{0}(f)=\{n\in P(f)\,|\,\forall m\geq 1(f^{m}(n)\geq n)\},$
i.e. the union of all cycles and the set of minimal representatives of cycles. By property (ii) there is an algorithm which for every $n\in\mathbb{N}$ verifies whether $n\in P(f)$, i.e. $P(f)$ is computable. Observe that $P_{0}(f)$ is computable too. Indeed, $1$ obviously belongs to $P_{0}(f)$. When $n\geq 2$ and $n\in P(f)$, then verifying if $f^{i}(n)\geq n$ for all $i\leq n$ we can check whether $n\in P_{0}(f)$ (apply (ii) again).
Since there is no component with two disjoint cycles we see that whenever $n,m\in P_{0}(f)$ and $n\neq m$, then $n$ and $m$ do not belong to the same connected component of the graph $\Gamma(f)$. Thus
@@
There is an algorithm which for every $n\in\mathbb{N}$ finds the $P_{0}(f)$-representative of the connected component of $n$. Indeed, if for example $n\notin P(f)$ then applying (iii) we can find $i\leq 2n$ such that $f^{i}(n)\in P(f)$ and later $j\leq n$ such that $f^{i+j}(n)\in P_{0}(f)$.
@@
  • $f(g(n,m))=g(n,m-1)$, and $f(h(n,m))=h(n,m-1)$.
We want $g,h$ to be computable functions. Since $P_{0}(f)$ and $P(f)$ are computable and the graph $\Gamma(f)$ is computable and $d$-regular the following rule gives required algorithm. Given $n\in P_{0}(f)$ and $m\geq 1$ and having defined $g(n,m-1)$ find the minimal $x$ such that $x\notin P(f)$ and $f(x)=g(n,m-1)$. Then let $g(n,m)=x$. The definition of $h(n,m)$ is similar. Clearly $g$ and $h$ are injective and have disjoint ranges.
@@
$f_{\star}(x)=\left\{\begin{array}[]{ll}g(y,m+2)&\text{ if }x=g(y,m)\text{ for }y\in P_{0}(f)\text{ and even }m\geq 0,\\ g(y,m-2)&\text{ if }x=g(y,m)\text{ for }y\in P_{0}(f)\text{ and odd }m\geq 3,\\ f^{2}(x)&\text{ if }x=h(y,m)\text{ for }y\in P_{0}(f)\text{ and }m\geq 2,\\ f(x)&\text{ otherwise.}\end{array}\right.$
We remind the reader that for each $n$ there is $i\leq 3n$ such that $f^{i}(n)\in P_{0}(f)$. Thus for any $m^{\prime}>3n$ the number $n$ does not belong to $\{g(y,m^{\prime}),h(y,m^{\prime})\}$. This guarantees that $f_{\star}$ is computable. Facts that $\Gamma(f_{\star})\in\mathcal{E}$, the function $f_{\star}$ does not have cycles and for each $x\in\mathbb{N}$ the size $|f^{-1}_{\star}(x)|$ is $d-1$ are proved in [Sch18, proof of Theorem 2.2]. Therefore the graph $\Gamma(f_{\star})$ is a computable $d$-regular forest.
To see the last statement of the theorem note that if $C\subset\mathbb{N}$ is a connected components of $\Gamma(f)$ then it is also the set of vertices of a connected component of $\Gamma(f_{\star})$. Since for each $n$ one can compute $n^{\prime}\in P_{0}(f)$ such that $n$ and $n^{\prime}$ are in same connected component of $\Gamma(f)$ we have an algorithm which for any $n$ and $m$ verifies whether $n$ and $m$ are in the same tree in $\Gamma(f_{\star})$. ∎
We now translate Theorem A into a statement about wobbling groups. The following theorem is Theorem B from the introduction. It is a computable version of Corollary 2.3 of [Sch18].
@@
Let the coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that $|E[F]|\geq 6|F|$ for all finite $F\subset\mathbb{N}$. Then there are two computable permutations $\sigma,\pi\in Sym(\mathbb{N})$ such that $\langle\sigma,\pi\rangle$ is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$.
@@
By Theorem A there is $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a computable $4$-regular forest. Moreover, there is an algorithm which for each $n,m\in\mathbb{N}$ recognizes if $n$ and $m$ are in the same connected component of $\Gamma(E^{\prime})$. Further, there is another algorithm which or each $n$ finds four natural numbers $n_{1}<n_{2}<n_{3}<n_{4}$ such that either $(n,n_{i})\in E^{\prime}$ or $(n_{i},n)\in E^{\prime}$. Since $\Gamma(E^{\prime})$ is a computable $4$-regular forest, for each $n,m\in\mathbb{N}$, we can compute $B_{m}(n)\subset\mathbb{N}$, the $m$-ball of $n$, by a uniform algorithm.
@@
The standard Cayley graph of the free group on two generators is isomorphic to the $4$-regular tree so it is clear that there exist permutations $\sigma$,$\pi$ as in the formulation. We will show that there exist computable ones. The construction is by induction. At every step we construct two finite partial isometries of $\Gamma(E^{\prime})$.
We will use the following notion. We say that a partial permutation $\sigma$ is loxodromic on a finite, connected set $D$, if:
  • it preserves distance i.e. for all $n,m\in D$ we have $d(n,m)=d(\sigma(n),\sigma(m))$;
  • there is a maximal line in the set $D\cup\sigma(D)$ i.e. there exists $d_{0},\ldots,d_{k}\in D\cup\sigma(D)$ for $k=\max\{d(a,b)|a,b\in D\}+1$, such that $\sigma(d_{0})=d_{1},\ldots,\sigma(d_{k-1})=d_{k}$ where $d_{0}\in D\setminus\sigma(D)$ and $d_{k}\in\sigma(D)\setminus D$.
At the first step of the construction of $\sigma$ and $\pi$ we find $1_{1},1_{2},1_{3},1_{4}$ and set $\sigma_{1}(1)=1_{1},\sigma_{1}(1_{2})=1,\pi_{1}(1)=1_{3},\pi_{1}(1_{4})=1$. Let $D_{1}=B_{1}(1)=\{1,1_{1},1_{2},1_{3},1_{4}\}$. We extend $\sigma_{1}$ and $\pi_{1}$ to loxodromic maps on $D_{1}$ such that $\sigma_{1}(D_{1})=B_{1}(\sigma_{1}(1))$ and $\pi_{1}(D_{1})=B_{1}(\pi_{1}(1))$.
@@
Assume that before the $n$-th step of the construction two finite partial permutations $\sigma_{n-1}$ and $\pi_{n-1})$ are defined so that they are loxodromic maps on some set $D_{n-1}$. For all $a\in D_{n-1}$ we set $\sigma_{n}(a)=\sigma_{n-1}(a)$ and $\pi_{n}(a)=\pi_{n-1}(a)$. Let $a_{n}$ be the first number in $\mathbb{N}\setminus D_{n-1}$. Let $X$ be the connected component of $\Gamma(E^{\prime})$ containing $a_{n}$. If it does not meet $D_{n-1}$ then we find the corresponding $(a_{n})_{1},(a_{n})_{2},(a_{n})_{3},(a_{n})_{4}$, and define $\sigma_{n}$ and $\pi_{n}$ on $\{a_{n},(a_{n})_{1},(a_{n})_{2},(a_{n})_{3},(a_{n})_{4}\}$ as at the first step. Then put $D_{n}=D_{n-1}\cup B_{1}(a_{n})$.
@@
It is worth mentioning a this stage that according the assumptions on $\Gamma(E^{\prime})$ the question if $X\cap D_{n-1}\not=\emptyset$ is decidable: just check if $a_{n}$ is in the same component with elements of $D_{n-1}$. Now assume that the answer is positive. Then find the minimal $i<n$ such that $a_{i}\in X$ and the minimal $m$ such that $a_{n}\in B_{m}(a_{i})$. Then we extend $\sigma_{n}$ and $\pi_{n}$ to a loxodromic partial permutations on $D_{n}=D_{n-1}\cup B_{m}(a_{i})$ in such way that for any $a\in D_{n}$ and any $i,j\in\{-1,1\}$ we have $\sigma^{i}_{n}(a)\neq\pi^{j}_{n}(a)$.
@@
Let $\sigma=\lim\limits_{n\in\mathbb{N}}\sigma_{n}$, and $\pi=\lim\limits_{n\in\mathbb{N}}\pi_{n}$. These permutations act on each connected component of $\Gamma(E^{\prime})$ by translations along some line, with lines corresponding to $\sigma$ and $\pi$ having one common vertex which is the first $n$ that belongs to component. Clearly there are no points which are fixed by non-identity of $\langle\sigma,\pi\rangle$ and this group is are a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$. Permutations $\sigma$ and $\pi$ are computable by the construction.

### v4

@@ 8. Proof of the main theorem
We remind the formulation of Theorem 4.1.
  • both $U$ and $V$ are identified with $\mathbb{N}\setminus\{0\}$
  • $E$ does not contain edges of the form $(u,v_{u})$
  • $\Gamma$ is fully reflected,
  • $\Gamma$ satisfies the $c.e.H.h.c.(d)$.
Then there exist a computable perfect (1,d1)(1,d-1)-matching of $\Gamma$, which realizes a computable (d1)(d-1) to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
@@
Let us apply the construction of Section 5. This construction works modulo Claims 5.2, 5.4, 5.6, 5.7. Claims 5.2 and 5.6 follow from Lemma 7.1. Claims 5.4 and 5.7 follow from Lemma 7.5. Since for every $n$ the union $\bigcup\limits_{i=0}^{n}M_{i}$ consists of (1,d1)(1,d-1)-fans, the final set of edges $M$ is a (1,d1)(1,d-1)-matching.
@@
For every $u\in U$ there is a step where an edge incident to $u$ is added to $M$. Furthermore, if the copy $v_{u}$ was not added to $M$ earlier, the construction ensures that $v_{u}$ is also added to $M$ in the second part of this step. This guarantees that $M$ is a perfect (1,d1)(1,d-1)-matching of the graph $\Gamma$. It also implies computability of this matching. Indeed, it is easy to see from the description given in Section 5 that a vertex $n$ appears in $M$ at the $n$-th step at latest. This follows from minimality of $u_{n}$ in $U^{(n-1)}$. Thus, for any $n$ all edges from $M$ of the forms $(n,j)$ and $(k,n)$ (the unique one) are already in $\bigcup\limits_{i=0}^{n}M_{i}$. Therefore, there is an algorithm which computes all such pairs from $M$ for any input $n$. We see that the function $f$ realized by this matching is computable.
Until the end of the proof we will abuse the notation and use a natural number $n$ both for vertices and numbers of steps.
@@
It remains to show that $f$ has controlled sizes of its cycle. To see condition $(i)$ of Definition 2.5 note that the edges $(u_{0},v^{0}_{0,1})$ and $(v_{u_{0}},u_{v^{0}_{0,1}})$ are added to $M$ at step 1, i.e. $f^{2}(u_{0})=f(u_{v^{0}_{0,1}})=u_{0}$. Since $u_{0}=1$, condition $(i)$ is satisfied.
@@
As a vertex $n$ appears in $M$ at the $n$-th step at latest, the length of a cycle created at the $n$-th step cannot be greater than $\max\{2,n\}$. In particular, if $n$ is in a cycle then $f^{i}(n)=n$ for some $i\leq n$ and condition $(ii)$ of Definition 2.5 is satisfied.
@@
It remains to show that condition $(iii)$ is satisfied. Let $\mathsf{f}_{i}$ be the partial function (living in $\mathbb{N}$) realized by $M_{i}$. Then $M_{i}=\{(\mathsf{f}_{i}(m),m):m\in\mathsf{Dom}(\mathsf{f}_{i})\}$ is the graph of $\mathsf{f}_{i}$. We denote by $f_{n}$ the partial function realized by $\bigcup\limits_{i=0}^{n}M_{i}$. Let $F_{n}=\{(f_{n}(m),m):m\in\mathsf{Dom}(f_{n})\}$ be the graph of $f_{n}$ on $\mathsf{Dom}(f_{n})\cup\mathsf{Rng}(f_{n})$. Then each $M_{i}$ with $i\leq n$, is a subgraph of $F_{n}$.
@@
Since each subgraph $M_{i}$ is connected, $F_{n}$ has at most $n+1$ connected components. Furthermore, one of the following properties holds:
  • $M_{n}$ is a connected component of $F_{n}$ and contains a cycle;
  • $M_{n}$ is connected and there is a vertex of degree $1$ in $F_{n-1}$ which is an image of a vertex from $M_{n}$.
@@
The first possibility arises in cases (2) and (3) of part 2 of the $(n+1)$-th step. The second possibility is the result of case (1) of this part of the step (see Section 5.10 and Remarks before part 2). Since each element is included into $M$ together with a fan, this property guarantees that $F_{n}$ consists only of vertices of degree $1$ and $d$. When a vertex has degree $1$, then its $F_{n}$-neighborhood is exactly its $f_{n}$-image. When a vertex has degree $d$, then its $f_{n}$-image and d1d-1 preimages are in $F_{n}$. In particular, each connected component of $F_{n}$ contains a cycle. The length of the cycle is not greater than $\max\{n,2\}$.
@@
Since the value $f_{n-1}(n)$ is defined, $n$ belongs to some connected component of $f_{n-1}$. Thus, there exist $k$ and $l$ such that $f_{n-1}^{k+l}(n)=f_{n-1}^{k}(n)$. These $k$ and $l$ work for the equality $f^{k+l}(n)=f^{k}(n)$. It remains to show that $k$ can be bounded by $2n$ and $l$ by $n$.
@@
The latter estimate is easy: the biggest cycle that can be constructed before the $n$-th step, is not longer than $\max\{n-1,2\}$. Below we will only use the inequality $l\leq n$ for simplicity.
@@
In order to show that $k$ is bounded by $2n$, let us estimate the size of a subset of $U$ that can be added to the matching $M$ in the process of the $n$-th step of the construction, in the situation when a new cycle is not created. It must consist of elements of $U^{(n-1)\perp}$ added to the matching at the iteration of part $2$ of the $n$-th step together with $u_{n}$. Therefore, we can bound it by the maximal possible number $|U^{(n-1)\perp}|+1$.
@@
Let us denote the number of elements from $U^{(s-1)\perp}$ added to $M$ at the $s$-th step by $\ell_{s}$. Then for each $m\in M_{s}$ we have $f^{\ell_{s}+1}(m)\in M_{j}$ where $j\leq s-1$. If no cycle is constructed at the $j$-th step then $f^{\ell_{j}+1}(f^{\ell_{s}+1}(m))\in M_{i}$ for some $i\leq j-1$. Iterating this argument and applying it to $m:=n$ we arrive at
@@
Since at the $n$-th step the value $|\bigcup\limits_{s=1}^{n-1}U^{(s)\perp}|$ does not exceed n1n-1,
@@ 9. Schneider’s theorem and computable entourages of coarse spaces
Let $d\geq 3$. Let $(\mathbb{N},\mathcal{E})$ be a non-amenable coarse space of a bounded geometry. Assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that for every finite $F\subseteq\mathbb{N}$ we have $|E[F]|\geq(d+2)|F|$.
Then there exists a computable $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a $d$-regular forest. Moreover, there exists an algorithm which for every $m,n\in\mathbb{N}$ recognizes if $m$ and $n$ are in the same connected component of $\Gamma(E^{\prime})$.
@@
Let $E$ be a highly computable symmetric entourage as in the formulation. Consider the graph $\Gamma(R)$ defined for the symmetric relation $R:=E\setminus\Delta_{\mathbb{N}}\subseteq\mathbb{N}\times\mathbb{N}$. Since the coarse space $(\mathbb{N},\mathcal{E})$ is of a bounded geometry, the neighborhood of each vertex of $\Gamma(R)$ is finite. Further, the graph $\Gamma(R)$ is highly computable by high computability of $E$.
Firstly, we apply Theorem 4.1 to $\Gamma(R)$. In order to do this, we only need to show that $\Gamma(R)$ satisfies $c.e.H.h.c.(d)$. To see the latter, note:
@@
In particular, the inequality $n\leq|X|$ implies $n\leq|N(X)|-d|X|\leq|N(X)|-\frac{1}{d}|X|$. Since the identity map on $\mathbb{N}$ is a total computable function, it follows that $\Gamma(R)$ satisfies $c.e.H.h.c.(d)$.
Now Theorem 4.1 provides a computable function $f$ that realizes a perfect (1,d1)(1,d-1) matching in the graph $\Gamma(R)$ and has controlled sizes of its cycles. This function will be used in the construction of $E^{\prime}\in\mathcal{E}$ as in the formulation of the theorem. We will adapt the proof of Theorem 2.2 of [schndr].
Since $f$ is a total, computable, surjective, and (d1)(d-1) to $1$ function, the graph of $f$ (denoted by $\Gamma(f)$) is computable and $d$-regular. We remind the reader that $f$ satisfies the following properties:
@@
[…]
i.e. $P(f)$ is the union of all cycles and $P_{0}(f)$ is the set of minimal representatives of cycles. By property (ii) there is an algorithm which for every $n\in\mathbb{N}$ verifies whether $n\in P(f)$, i.e. $P(f)$ is computable. Observe that $P_{0}(f)$ is computable too. Indeed, the number $1$ obviously belongs to $P_{0}(f)$. When $n\geq 2$ and $n\in P(f)$, then verifying if $f^{i}(n)\geq n$ for all $i\leq n$ we can check whether $n\in P_{0}(f)$ (apply (ii) again).
Since there is no component with two disjoint cycles, we see that whenever $n,m\in P_{0}(f)$ and $n\neq m$, then $n$ and $m$ do not belong to the same connected component of the graph $\Gamma(f)$. Thus,
@@
Further, there is an algorithm which for every $n\in\mathbb{N}$ finds the $P_{0}(f)$-representative of the connected component of $n$. Indeed, if for example $n\notin P(f)$, then applying (iii) we compute $i\leq 2n$ such that $f^{i}(n)\in P(f)$, and later we find $j\leq n$ such that $f^{i+j}(n)\in P_{0}(f)$.
@@
  • f(g(n,m))=g(n,m1)f(g(n,m))=g(n,m-1), and f(h(n,m))=h(n,m1)f(h(n,m))=h(n,m-1).
We want $g,h$ to be computable functions. Since $P_{0}(f)$ and $P(f)$ are computable and the graph $\Gamma(f)$ is computable and $d$-regular, the following rule gives the required algorithm. Given $n\in P_{0}(f)$ and $m\geq 1$ and having already defined g(n,m1)g(n,m-1), find the minimal $x$ such that $x\notin P(f)$ and f(x)=g(n,m1)f(x)=g(n,m-1). Then let $g(n,m)=x$. The definition of $h(n,m)$ is similar. Clearly $g$ and $h$ are injective and have disjoint ranges.
@@
[…]
We remind the reader that for each $n$ there is $i\leq 3n$ such that $f^{i}(n)\in P_{0}(f)$. Thus for any $m^{\prime}>3n$ the number $n$ does not belong to $\{g(y,m^{\prime}),h(y,m^{\prime})\}$. This guarantees that $f_{\star}$ is computable. The facts that $\Gamma(f_{\star})\in\mathcal{E}$, the function $f_{\star}$ does not have cycles, and for each $x\in\mathbb{N}$ the size $|f^{-1}_{\star}(x)|$ is d1d-1 are proved in [schndr, proof of Theorem 2.2]. Therefore, the graph $\Gamma(f_{\star})$ is a computable $d$-regular forest.
To see the last statement of the theorem note that if $C\subset\mathbb{N}$ is a connected components of $\Gamma(f)$ then it is also the set of vertices of a connected component of $\Gamma(f_{\star})$. Since for each $n$ one can compute $n^{\prime}\in P_{0}(f)$ such that $n$ and $n^{\prime}$ are in same connected component of $\Gamma(f)$, there is a straightforward algorithm which for any $n$ and $m$ verifies whether $n$ and $m$ are in the same tree in $\Gamma(f_{\star})$. ∎
We now translate Theorem A into a statement about wobbling groups. The following theorem is Theorem B from the introduction. It is a computable version of Corollary 2.3 of [schndr].
@@
Let $(\mathbb{N},\mathcal{E})$ be a non-amenable coarse space of a bounded geometry. Assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that $|E[F]|\geq 6|F|$ for all finite $F\subset\mathbb{N}$. Then there are two computable permutations $\sigma,\pi\in Sym(\mathbb{N})$ such that $\langle\sigma,\pi\rangle$ is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$.
@@
By Theorem A there is $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a computable $4$-regular forest. Moreover, there is an algorithm which for each $n,m\in\mathbb{N}$ recognizes if $n$ and $m$ are in the same connected component of $\Gamma(E^{\prime})$. Since $\Gamma(E^{\prime})$ is a computable $4$-regular forest, there is another algorithm which for each $n$ finds four natural numbers $n_{1}<n_{2}<n_{3}<n_{4}$ such that for every $n_{i}$ either $(n,n_{i})\in E^{\prime}$ or $(n_{i},n)\in E^{\prime}$. Furthermore, for each input $n,m\in\mathbb{N}$, the output $B_{m}(n)\subset\mathbb{N}$, the $m$-ball of $n$, can be effectively found by a uniform algorithm.
@@
The standard Cayley graph of the free two generated group is isomorphic to the $4$-regular tree, so it is clear that there exist permutations $\sigma$$\pi$ as in the formulation. We will show that there exist computable ones. The construction is by induction. At every step we construct two finite partial permutations of $\Gamma(E^{\prime})$.
At the first step of the construction of $\sigma$ and $\pi$ put $D_{1}=B_{1}(1)=\{1,1_{1},1_{2},1_{3},1_{4}\}$ and define $\sigma_{1}$ and $\pi_{1}$ on $D_{1}$ so that $\sigma_{1}(1)=1_{1},\sigma_{1}(1_{2})=1,\pi_{1}(1)=1_{3},\pi_{1}(1_{4})=1$, and for every $i\leq 4$, $\{\sigma_{1}(1_{i})^{\pm 1},\pi_{1}(1_{i})^{\pm 1}\}\subset B_{1}(1_{i})$ with $\sigma^{\pm 1}_{1}(1_{i})\not=\pi^{\pm 1}_{1}(1_{i})$ (whenever the corresponding preimages are defined); the set $D_{1}\cup\sigma_{1}(D_{1})\cup\pi_{1}(D_{1})$ consists of 11 elements.
Assume that before the $n$-th step of the construction two finite partial permutations $\sigma_{n-1}$ and $\pi_{n-1}$ are defined on some set $D_{n-1}$ such that for all $a\in D_{n-1}$, $\{\sigma^{\pm 1}_{n-1}(a),\pi^{\pm 1}_{n-1}(a)\}\subset B_{1}(a)$.
@@
At step $n$ we first put $\sigma_{n}(a)=\sigma_{n-1}(a)$ and $\pi_{n}(a)=\pi_{n-1}(a)$ for all $a\in D_{n-1}$. Let $a_{n}$ be the first number in $\mathbb{N}\setminus D_{n-1}$. Let $X$ be the connected component of $\Gamma(E^{\prime})$ containing $a_{n}$. If it does not meet $D_{n-1}$, then we find the corresponding $(a_{n})_{1},(a_{n})_{2},(a_{n})_{3},(a_{n})_{4}$, and define $\sigma_{n}$ and $\pi_{n}$ on $\{a_{n},(a_{n})_{1},(a_{n})_{2},(a_{n})_{3},(a_{n})_{4}\}$ exactly as at the first step. Then put $D_{n}=D_{n-1}\cup B_{1}(a_{n})$.
@@
It is worth mentioning at this stage that according to the assumptions on $\Gamma(E^{\prime})$, the question if $X\cap D_{n-1}\not$$\emptyset$ is decidable: just check if $a_{n}$ is in the same component with elements of $D_{n-1}$.
Now, assume that the answer is positive. Then find the minimal $i<n$ such that $a_{i}\in X$ and the minimal $m$ such that $a_{n}\in B_{m}(a_{i})$. Put $D_{n}=D_{n-1}\cup B_{m}(a_{i})$. Extend $\sigma_{n-1}\upharpoonright_{X}$ and $\pi_{n-1}\upharpoonright_{X}$ to partial permutations defined on $B_{m}(a_{i})$ such that for every $a\in B_{m}(a_{i})$, $\{\sigma^{\pm 1}_{n}(a),\pi^{\pm 1}_{n}(a)\}\subset B_{1}(a)$, and $\sigma^{\pm 1}_{n}(a)\neq\pi^{\pm 1}_{n}(a)$ (whenever the corresponding preimages are defined).
The set $B_{m}(a)$ can be viewed as the set of all $\mathsf{w}(a)$ where $\mathsf{w}$ is a group word of length $\leq n$ of the alphabet $\{\ell,r\}$. Then $\sigma_{n}$ (resp. $\pi_{n}$) takes $\mathsf{w}(a)$ with the image undefined before, to $\ell\mathsf{w}(a)$ (resp. $r\mathsf{w}(a))$.
@@
Let $\sigma=\lim\limits_{n\in\mathbb{N}}\sigma_{n}$, and $\pi=\lim\limits_{n\in\mathbb{N}}\pi_{n}$. These permutations act on each connected component of $\Gamma(E^{\prime})$ by mapping elements to their neighbors. Clearly, there are no points which are fixed by non-identity elements of the group $\langle\sigma,\pi\rangle$. It is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$. Permutations $\sigma$ and $\pi$ are computable by the construction. ∎